In an experiment, a 10.67-g sample of magnesium was heated in a stream of pure oxygen, and was found to increase in mass by 1.004-g. Calculate the empirical formula of magnesium oxide.


1
Expert's answer
2021-01-27T04:40:54-0500

Assuming that the values given are okay;

Mass of Mg = 10.67g

Moles of Mg "=\dfrac{10.67g}{24.305g/mol} = 0.4378mol"



Mass increase = Mass of oxygen

Thus mass of O = 1.004g

Moles of O atoms "=\dfrac{1,004g}{15.999g/mol} = 0.0628mol"


Mole ratio of Mg : O = 0.4378 : 0.0628

= "\dfrac{0.4378}{0.0628}: \dfrac{0.0628}{0.0628} = 6.9:1=7:1"


The empirical formula is Mg7O